# Robust parameter estimation within minutes on gravitational wave signals from binary neutron star inspirals

### v1

@@ Abstract
The gravitational waves emitted by binary neutron star inspirals contain information on nuclear matter above saturation density. However, extracting this information and conducting parameter estimation remains a computationally challenging and expensive task. Wong et al. introduced Jim Wong et al. (2023a), a parameter estimation pipeline that combines relative binning and jax features such as hardware acceleration and automatic differentiation into a normalizing flow-enhanced sampler for gravitational waves from binary black hole (BBH) mergers. In this work, we extend the Jim framework to analyze gravitational wave signals from binary neutron star (BNS) mergers with tidal effects included. We demonstrate that Jim can be used for full Bayesian parameter estimation of gravitational waves from BNS mergers within a few tens of minutes, which includes the training of the normalizing flow and computing the reference parameters for relative binning. For instance, Jim can analyze GW170817 in $26$ minutes ($33$ minutes) of total wall time using the TaylorF2 (IMRPhenomD_NRTidalv2) waveform, and GW190425 in around $21$ minutes for both waveforms. We highlight the importance of such an efficient parameter estimation pipeline for several science cases as well as its ecologically friendly implementation of gravitational wave parameter estimation.
@@ I Introduction
NS are the remnants of core-collapse supernovae and consist of the densest matter ever observed in the Universe Lattimer (2012); Özel and Freire (2016). With densities up to a few times the nuclear saturation density $n_{\rm sat}=2.7\times 10^{14}{\rm g\ cm}^{-3}$, neutron stars are the perfect laboratories for studying the properties and behavior of ultra-dense matter. The equation of state (EOS) relates the pressure, temperature, and energy density within the interior of NSs and is not completely understood Tews et al. (2022). It is defined by the fundamental degrees of freedom within the NS and the interactions among them. Each proposed EOS uniquely dictates the global structure of NSs, influencing their masses, radii, and tidal deformabilities Hinderer et al. (2010); Damour and Nagar (2009); Lindblom (1992). Thus, astronomical observations of NSs allow one to constrain the EOS in NSs Flanagan and Hinderer (2008); Del Pozzo et al. (2013); Agathos et al. (2015); Demorest et al. (2010); Antoniadis et al. (2013); Arzoumanian et al. (2018); Cromartie et al. (2019); Miller et al. (2019); Riley et al. (2019); Miller et al. (2021); Riley et al. (2021).
@@
Since the first detection of gravitational waves from a BNS merger in 2017 Abbott et al. (2017), by Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015), GW astronomy has become an important channel for astronomical observation of NSs. An essential step for extracting physics from the GW data is performing parameter estimation (PE) on it.
@@
Within the community, multiple central processing unit (CPU)-based PE software packages have been developed, which include LALInference Veitch et al. (2015), PyCBC Inference Biwer et al. (2019) and Bilby Ashton et al. (2019); Romero-Shaw et al. (2020); Smith et al. (2020). These packages are robust and have been used for analyzing multiple GW events Abbott et al. (2017, 2020). Yet, they are also known for being computationally expensive and having a large carbon footprint, especially for analyzing BNS events. For these reasons, a lot of effort has been put into developing methods to speed up PE, e.g., relative binning Cornish (2010); Zackay et al. (2018) and reduced order quadrature (ROQCanizares et al. (2013, 2015), which show substantial improvements Krishna et al. (2023); Morisaki et al. (2023).
@@
In addition, there have been proposals to accelerate PE through the use of graphical processing unit (GPU)-based software or machine learning (ML) techniques. One such technique, known as normalizing flows, has been gaining popularity in the field. For instance, Dingo uses normalizing flows pretrained on simulated data to approximate the posterior distribution Dax et al. (2021, 2023); Bhardwaj et al. (2023). Similarly, normalizing flows have been used to accelerate nested sampling for GW data analysis by nessai Williams et al. (2021, 2023). Recently, a GPU-based PE software, Jim Wong et al. (2023a), has been introduced, which accelerates PE with Markov chain Monte Carlo (MCMC) samplers using normalizing flows. In Ref. Wong et al. (2023a), the authors demonstrated that Jim can conduct PE on GW signals within minutes of total wall time on a single GPU.
@@
In this paper, we extend the capabilities of Jim to analyze GWs from BNS mergers, in particular to also infer the tidal deformabilities of GW signals. The updated Jim framework is crucial for low-latency PE, necessary for efficient telescopes’ responses on BNS detections, for handling the PE for a large number of BNS events, and to aid the multi-messenger analysis on BNS via integration with a nuclear physics multi-messenger framework, e.g., NMMA Pang et al. (2023). As we show in more detail below, Jim performs these efficient runs at a lower carbon footprint compared to other existing pipelines.
@@
This paper is structured as follows. In Sec. II, we give an overview of Jim and our methods, with their validation shown in Sec. III. In Sec. IV, we apply our methods on the two BNS GW detections to date, namely GW170817 Abbott et al. (2017) and GW190425 Abbott et al. (2020). We compare our work to other methods and discuss implications for future work in Sec. V. Concluding remarks and future perspectives are provided in Sec. VI.
@@ II.1 Parameter estimation
Based on Bayes’ theorem, the posterior distribution of the source parameters $\bm{\theta}$ of a GW signal $d$, denoted by $p(\bm{\theta}|d)$, is given by Veitch and Vecchio (2010)
@@
with $S_{n}(f)$ being the one-sided power spectral density (PSD), $\tilde{x}(f)$ being the Fourier transform of $x(t)$, and the asterisk denoting the complex conjugate. In this work, the frequency range of the integral is chosen such that it covers the frequency span of the signal, which for BNS signals is usually taken as $[20,2048]$ Hz. A complete list of parameters and corresponding definitions can be found in Ref. Romero-Shaw et al. (2020). The notation used in this work can be found in Tab. 2.
@@ II.2 Waveform approximants
Several families of waveform approximants have been developed and are used in GW data analysis. These families are mainly divided into the post-Newtonian (PN) waveforms Buonanno et al. (2009); Blanchet (2014), the effective one-body waveforms Buonanno and Damour (1999, 2000); Damour and Nagar (2011), such as TEOBResumS Bernuzzi et al. (2015); Nagar et al. (2018); Akcay et al. (2019) or SEOBNRv4T Hinderer et al. (2016); Steinhoff et al. (2016) for BNS inspirals, the inspiral-merger-ringdown phenomenological (IMRPhenom) waveforms, such as IMRPhenomD for BBH inspirals Husa et al. (2016); Khan et al. (2016), or surrogate models based on numerical-relativity simulations Varma et al. (2019). In this work, we will limit ourselves to the frequency domain approximants TaylorF2, which is a PN approximant, and IMRPhenomD_NRTidalv2 Dietrich et al. (2017, 2019), which are both frequently used in the analysis of BNS events. We do not consider the recently introduced IMRPhenomXAS_NRTidalv2 Colleoni et al. (2023) or NRTidalv3 models Abac et al. (2024), since these waveform models are not released at the time of writing.
@@ II.3 Relative binning
Relative binning Zackay et al. (2018); Cornish (2021); Krishna et al. (2023) is a technique to speed up the evaluation of the likelihood function given in Eq. (2). Given a reference parameter $\bm{\theta}_{\rm ref}$, the ratio $r(f)$ of the waveform of an arbitrary parameter $\bm{\theta}$ against the reference waveform is given by
@@
$r(f)\approx\begin{cases}r_{0}(b_{1})+r_{1}(b_{1})(f-f_{\rm m}(b_{1})),&f\in b_{1}\\ r_{0}(b_{2})+r_{1}(b_{2})(f-f_{\rm m}(b_{2})),&f\in b_{2}\\ \qquad\qquad\qquad\vdots&\\ r_{0}(b_{n})+r_{1}(b_{n})(f-f_{\rm m}(b_{n})).&f\in b_{n}\\ \end{cases}$ (5)
@@
The bin placement often follows the approach described in Ref. Zackay et al. (2018). In this scheme, one assumes that the variations in the ratio of amplitudes can be neglected. The error on the ratio is therefore determined by the deviations in the phase, and one can place the bin edges in such a way to ensure that the deviation in the phase across a bin is below a certain desired threshold. This deviation is approximated with an ansatz of the phase based on PN theory to remove the dependence on the waveform’s parameters. We refer readers to Refs. Zackay et al. (2018); Krishna et al. (2023) for further details.
@@ II.3.1 Caveats for applying relative binning with NRTidal-family waveforms
The above assumption regarding the amplitude ratio holds for a majority of waveform models (e.g., TaylorF2). However, it is invalid for the IMRPhenomD_NRTidalv2 waveform model. This waveform model modifies its underlying point-particle baseline, IMRPhenomD, with closed-form expressions for the tidal contributions to the phase and amplitude. These are modeled by Padé approximants, such that the PN ansatz for the phase used in placing the bin edges is no longer a valid approximation. Moreover, IMRPhenomD_NRTidalv2 additionally tapers the waveform so that its amplitude transitions smoothly to zero in the range ${[f_{\rm merger},1.2f_{\rm merger}]}$, where the merger frequency $f_{\rm merger}$ depends on the intrinsic source parameters and was estimated through a phenomenological relation Dietrich et al. (2019). For larger deviations from the reference parameters, this causes the ratio of amplitudes in Eq. (4) to vary significantly, causing a breakdown of the relative binning scheme. Moreover, we observed that, for some parameter regimes for which the reference waveform has a low merger frequency, the ratio of amplitudes becomes numerically unstable. Therefore, we generate the reference waveform without the tapering window when analyzing simulated signals in Sec.III below.
@@ II.3.2 Reference parameters
Relative binning requires a set of reference parameters as input that, ideally, lie close to the best-fit parameters. When analyzing simulated signals, we set the reference parameters equal to the injected parameters. For analyzing real events, we first perform a heuristic search to obtain an estimate of the maximum likelihood parameters. More specifically, we employ the covariance matrix adaptation evolution strategy (CMA-ES) Hansen and Ostermeier (2001) as implemented in evosax Lange (2022) before running Jim to obtain the reference parameters used in the PE run.
@@ II.4 Jim
Jim Wong et al. (2023a) is a PE pipeline to analyze GW events involving BBH mergers that is able to run on CPUs, GPUs or tensor processing units (TPUs). In Ref. Wong et al. (2023a), the authors demonstrated Jim’s capability of performing PE in the order of minutes of total wall-time, including training of the normalizing flow. In this work, we extend Jim’s capability of performing PE on GW signals from BNS inspirals with tidal effects included.
@@
Jim is implemented in jax Frostig et al. (2018), a high-performance numerical computation library. jax has a number of desirable features for GW data analysis, such as i) automatic differentiation, allowing the use of gradient-based MCMC samplers, ii) native support for hardware accelerators, such as GPUs or TPUs, and iii) just-in-time (JIT) compilation to further accelerate the execution of the code.
@@
For our waveform generator, we make use of the ripple Edwards et al. (2023), a jax package for differentiable waveform approximants. To analyze BNS signals, we extend ripple with the TaylorF2 and IMRPhenomD_NRTidalv2 waveform models. [^1] While TaylorF2 and IMRPhenomD_NRTidalv2 already have an existing jax implementation in gwfast Iacovelli et al. (2022), we choose to provide an independent implementation in ripple in order to interface more easily with Jim and flowMC. [^2] We note that Ref. Chia et al. (2023) has used TaylorF2 with Jim for PE. However, here we provide an extensive validation of the implementation and provide detailed benchmarks.
@@
For the MCMC sampler, we employ flowMC, a jax-based MCMC sampler that makes use of gradient-based samplers and is enhanced by normalizing flows Wong et al. (2023b); Gabrié et al. (2022). The principle behind the flowMC sampler is summarized by the diagram in Fig. 1. flowMC combines a local and global proposal distribution to improve the sampler’s efficiency. For the local sampler, we use the Metropolis-adjusted Langevin algorithm (MALA) routine Grenander and Miller (1994), which exploits the gradient of the posterior distribution to evolve the Markov chains. The global sampler is parameterized by a normalizing flow (NF), which is implemented with equinox Kidger and Garcia (2021). NFs are deep generative models that offer tractable approximations of complex probability distributions, allowing for efficient and precise sampling as well as density evaluation. In flowMC, the NF is trained from the samples generated by the local sampler and subsequently used as proposal distribution for the global sampling phase.
@@
The sampler settings can be either optimized for speed or accuracy. In this work, we use settings that focus on the latter. For instance, to improve the robustness of the sampler, we use a stopping threshold. Once the NF achieves a preset threshold value for the mean acceptance rate during the Metropolis-Hastings steps, the NF is frozen, and the final production samples are produced by only running the local and global samplers (known as the production loop) for a fixed amount of epochs. Training of the NF dominates the wall time, while the production loop only takes $1$-$2$ minutes.

### v2

@@ Abstract
The gravitational waves emitted by binary neutron star inspirals contain information on nuclear matter above saturation density. However, extracting this information and conducting parameter estimation remains a computationally challenging and expensive task. Wong et al. introduced Jim Wong:2023lgb, a parameter estimation pipeline that combines relative binning and jax features such as hardware acceleration and automatic differentiation into a normalizing flow-enhanced sampler for gravitational waves from binary black hole (BBH) mergers. In this work, we extend the Jim framework to analyze gravitational wave signals from binary neutron star (BNS) mergers with tidal effects included. We demonstrate that Jim can be used for full Bayesian parameter estimation of gravitational waves from BNS mergers within a few tens of minutes, which includes the training of the normalizing flow and computing the reference parameters for relative binning. For instance, Jim can analyze GW170817 in $20$ minutes ($31$ minutes) of total wall time using the TaylorF2 (IMRPhenomD_NRTidalv2) waveform, and GW190425 in around $21$ minutes ($25$ minutes). We highlight the importance of such an efficient parameter estimation pipeline for several science cases as well as its ecologically friendly implementation of gravitational wave parameter estimation.
@@ I Introduction
NS are the remnants of core-collapse supernovae and consist of the densest matter ever observed in the Universe Lattimer:2012nd; Ozel:2016oaf. With densities up to a few times the nuclear saturation density $n_{\rm sat}=2.7\times 10^{14}{\rm g\ cm}^{-3}$, neutron stars are the perfect laboratories for studying the properties and behavior of ultra-dense matter. The equation of state (EOS) relates the pressure, temperature, and energy density within the interior of NSs and is not completely understood Ozel:2016oaf. It is defined by the fundamental degrees of freedom within the NS and the interactions among them. Each proposed EOS uniquely dictates the global structure of NSs, influencing their masses, radii, and tidal deformabilities Hinderer:2009ca; Damour:2009vw; Lindblom1992-zb. Thus, astronomical observations of NSs allow one to constrain the EOS in NSs Flanagan:2007ix; DelPozzo:2013ala; Agathos:2015uaa; Demorest:2010bx; Antoniadis:2013pzd; Arzoumanian:2017puf; Cromartie:2019kug; Miller:2019cac; Riley:2019yda; Miller:2021qha; Riley:2021pdl.
@@
Since the first detection of gravitational waves from a BNS merger in 2017 LIGOScientific:2017vwq, by Advanced LIGO LIGOScientific:2014pky and Advanced Virgo VIRGO:2014yos, GW astronomy has become an important channel for astronomical observation of NSs. An essential step for extracting physics from the GW data is performing parameter estimation (PE) on it.
@@
Within the community, multiple central processing unit (CPU)-based PE software packages have been developed, which include LALInference Veitch:2014wba, PyCBC Inference Biwer:2018osg and Bilby Ashton:2018jfp; Romero-Shaw:2020owr; Smith:2019ucc. These packages are robust and have been used for analyzing multiple GW events LIGOScientific:2017vwq; LIGOScientific:2020aai. Yet, they are also known for being computationally expensive and having a large carbon footprint, especially for analyzing BNS events. For these reasons, a lot of effort has been put into developing methods to speed up PE, e.g., relative binning Cornish:2010kf; Zackay:2018qdy and reduced order quadrature (ROQCanizares:2013ywa; Canizares:2014fya, which show substantial improvements Krishna:2023bug; Morisaki:2023kuq.
@@
In addition, there have been proposals to accelerate PE through the use of graphical processing unit (GPU)-based software or machine learning (ML) techniques. One such technique, known as normalizing flows, has been gaining popularity in the field. For instance, Dingo uses normalizing flows pretrained on simulated data to approximate the posterior distribution Dax:2021tsq; Dax:2022pxd; Bhardwaj:2023xph. Similarly, normalizing flows have been used to accelerate nested sampling for GW data analysis by nessai Williams:2021qyt; Williams:2023ppp. Recently, a GPU-based PE software, Jim Wong:2023lgb, has been introduced, which accelerates PE with Markov chain Monte Carlo (MCMC) samplers using normalizing flows. In Ref. Wong:2023lgb, the authors demonstrated that Jim can conduct PE on GW signals within minutes of total wall time on a single GPU.
@@
In this paper, we extend the capabilities of Jim to analyze GWs from BNS mergers, in particular to also infer the tidal deformabilities of GW signals. The updated Jim framework is crucial for low-latency PE, necessary for efficient telescopes’ responses on BNS detections, for handling the PE for a large number of BNS events, and to aid the multi-messenger analysis on BNS via integration with a nuclear physics multi-messenger framework, e.g., NMMA Pang:2022rzc. As we show in more detail below, Jim performs these efficient runs at a lower carbon footprint compared to other existing pipelines.
@@
This paper is structured as follows. In Sec. II, we give an overview of Jim and our methods, with their validation shown in Sec. III. In Sec. IV, we apply our methods on the two BNS GW signals detected to date, namely GW170817 LIGOScientific:2017vwq and GW190425 LIGOScientific:2020aai. We compare our work to other methods and discuss implications for future work in Sec. V. Concluding remarks and future perspectives are provided in Sec. VI.
@@ II.1 Parameter estimation
Based on Bayes’ theorem, the posterior distribution of the source parameters $\bm{\theta}$ of a GW signal $d$, denoted by $p(\bm{\theta}|d)$, is given by Veitch:2009hd
@@
with $S_{n}(f)$ being the one-sided power spectral density (PSD), $\tilde{x}(f)$ being the Fourier transform of $x(t)$, and the asterisk denoting the complex conjugate. In this work, the frequency range of the integral is chosen such that it covers the frequency span of the signal, which for BNS signals is usually taken as $[20,2048]$ Hz. A complete list of parameters and corresponding definitions can be found in Ref. Romero-Shaw:2020owr. The notation used in this work can be found in Tab. 2.
@@ II.2 Waveform approximants
Several families of waveform approximants have been developed and are used in GW data analysis. These families are mainly divided into the post-Newtonian (PN) waveforms Buonanno:2009zt; Blanchet:2013haa, the effective one-body waveforms Buonanno:1998gg; Buonanno:2000ef; Damour:2009zoi, such as TEOBResumS Bernuzzi:2014owa; Nagar:2018zoe; Akcay:2018yyh or SEOBNRv4T Hinderer:2016eia; Steinhoff:2016rfi for BNS inspirals, the inspiral-merger-ringdown phenomenological (IMRPhenom) waveforms, such as IMRPhenomD for BBH inspirals Husa:2015iqa; Khan:2015jqa, or surrogate models based on numerical-relativity simulations Varma:2019csw. In this work, we will limit ourselves to the frequency domain approximants TaylorF2, which is a PN approximant, and IMRPhenomD_NRTidalv2 Dietrich:2017aum; Dietrich:2019kaq, which are both frequently used in the analysis of BNS events. We do not consider the recently introduced IMRPhenomXAS_NRTidalv2 Colleoni:2023czp or NRTidalv3 models Abac:2023ujg, since these waveform models are not released at the time of writing.
@@ II.3 Relative binning
Relative binning Zackay:2018qdy; Cornish:2021lje; Krishna:2023bug is a technique to speed up the evaluation of the likelihood function given in Eq. (2). Given a reference parameter $\bm{\theta}_{\rm ref}$, the ratio $r(f)$ of the waveform of an arbitrary parameter $\bm{\theta}$ against the reference waveform is given by
@@
[…]
@@
The bin placement often follows the approach described in Ref. Zackay:2018qdy. In this scheme, one assumes that the variations in the ratio of amplitudes can be neglected. The error on the ratio is therefore determined by the deviations in the phase, and one can place the bin edges in such a way to ensure that the deviation in the phase across a bin is below a certain desired threshold. This deviation is approximated with an ansatz of the phase based on PN theory to remove the dependence on the waveform’s parameters. We refer readers to Refs. Zackay:2018qdy; Krishna:2023bug for further details.
@@ II.3.1 Caveats for applying relative binning with NRTidal-family waveforms
The above assumption regarding the amplitude ratio holds for a majority of waveform models (e.g., TaylorF2). However, it is invalid for the IMRPhenomD_NRTidalv2 waveform model. This waveform model modifies its underlying point-particle baseline, IMRPhenomD, with closed-form expressions for the tidal contributions to the phase and amplitude. These are modeled by Padé approximants, such that the PN ansatz for the phase used in placing the bin edges is no longer a valid approximation. Moreover, IMRPhenomD_NRTidalv2 additionally applies a Planck taper McKechan:2010kp to the waveform so that its amplitude transitions smoothly to zero in the range $f_{\rm merger},1.2f_{\rm merger}]$}, where the merger frequency $f_{\rm merger}$ depends on the intrinsic source parameters and was estimated through a phenomenological relation Dietrich:2019kaq. For larger deviations from the reference parameters, this causes the ratio of amplitudes in Eq. (4) to vary significantly, causing a breakdown of the relative binning scheme. Moreover, we observed that, for some parameter regimes for which the reference waveform has a low merger frequency, the ratio of amplitudes becomes numerically unstable. Therefore, Jim generates the reference waveform without the tapering window when analyzing signals with the IMRPhenomD_NRTidalv2 waveform model, whereas we still include the taper in the sampled waveforms. The injected signals of our injections in Sec. III also apply the taper. We note that bilby, contrary to Jim, includes the taper in the reference waveform.
@@ II.3.2 Reference parameters
Relative binning requires a set of reference parameters as input that, ideally, lie close to the best-fit parameters. When analyzing simulated signals, we set the reference parameters equal to the injected parameters. For analyzing real events, we first perform a heuristic search to obtain an estimate of the maximum likelihood parameters. More specifically, we employ the covariance matrix adaptation evolution strategy (CMA-ES) CMA_ES_reference_paper as implemented in evosax lange2022evosax before running Jim to obtain the reference parameters used in the PE run. When searching for a set of reference parameters with the IMRPhenomD_NRTidalv2 waveform, we include its amplitude taper.
@@ II.4 Jim
Jim Wong:2023lgb is a PE pipeline to analyze GW events involving BBH mergers that is able to run on CPUs, GPUs or tensor processing units (TPUs). In Ref. Wong:2023lgb, the authors demonstrated Jim’s capability of performing PE in the order of minutes of total wall-time, including training of the normalizing flow. In this work, we extend Jim’s capability of performing PE on GW signals from BNS inspirals with tidal effects included.
@@
Jim is implemented in jax frostig2018compiling, a high-performance numerical computation library. jax has a number of desirable features for GW data analysis, such as i) automatic differentiation, allowing the use of, e.g., gradient-based MCMC samplers, ii) native support for hardware accelerators, such as GPUs or TPUs, and iii) just-in-time (JIT) compilation to further accelerate the execution of the code.
@@
For our waveform generator, we make use of ripple Edwards:2023sak, a jax package for differentiable waveform approximants. To analyze BNS signals, we extend ripple with the TaylorF2 and IMRPhenomD_NRTidalv2 waveform models. [^1] While TaylorF2 and IMRPhenomD_NRTidalv2 already have an existing jax implementation in gwfast Iacovelli:2022mbg, we choose to provide an independent implementation in ripple in order to interface more easily with Jim and flowMC. [^2] We note that Ref. Chia:2023tle has used TaylorF2 with Jim for PE. However, here we provide an extensive validation of the implementation and provide detailed benchmarks.
@@
For the MCMC sampler, we employ flowMC, a jax-based MCMC sampler that makes use of gradient-based samplers and is enhanced by normalizing flows Wong:2022xvh; Gabrie:2021tlu. The flowMC sampler is summarized by the diagram in Fig. 1. flowMC combines a local and global proposal distribution to improve the sampler’s efficiency. For the local sampler, we use the Metropolis-adjusted Langevin algorithm (MALA) routine grenander1994representations, which exploits the gradient of the posterior distribution to evolve the Markov chains. The global sampler is parameterized by a normalizing flow (NF), which is implemented with equinox kidger2021equinox. NFs are deep generative models that offer tractable approximations of complex probability distributions, allowing for efficient and precise sampling as well as density evaluation. In flowMC, the NF is trained from the samples generated by the local sampler and subsequently used as proposal distribution for the global sampling phase.
@@
The sampler settings can be either optimized for speed or accuracy. In this work, we use settings that focus on the latter. For instance, to improve the robustness of the sampler, we use a stopping threshold. Once the NF achieves a preset threshold value for the mean acceptance rate during the Metropolis-Hastings steps, the NF is frozen, and the final production samples are produced by only running the local and global samplers (known as the production loop) for a fixed amount of epochs. Training of the NF dominates the wall time, while the production loop only takes 11-22 minutes.